# ActSonic: Recognizing Everyday Activities from Inaudible Acoustic Waves Around the Body

### v2

@@ Abstract.
We present ActSonic, an intelligent, low-power active acoustic sensing system integrated into eyeglasses that can recognize $27$ different everyday activities (e.g., eating, drinking, toothbrushing) from inaudible acoustic waves around the body with a time resolution of one second. It only needs a pair of miniature speakers and microphones mounted on each hinge of eyeglasses to emit ultrasonic waves to create an acoustic aura around the body. Based on the position and motion of various body parts, the acoustic signals are reflected with unique patterns captured by the microphone and analyzed by a customized self-supervised deep learning framework to infer the performed activities. ActSonic was deployed in a user study with 19 participants across 19 households to evaluate its efficacy. Without requiring any training data from a new user (leave-one-participant-out evaluation), ActSonic was able to detect 27 activities, achieving an average F1-score of 86.6% in fully unconstrained scenarios and 93.4% in prompted settings at participants’ homes.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: IMWUT
@@ 1. Introduction
Acquiring precise body pose data from various body parts is essential for accurately distinguishing different human activities. However, achieving this with everyday wearable devices (e.g., eyeglasses) has presented persistent challenges: Wearable cameras (Price et al., 2022; Liu et al., 2023) consume significant power, rapidly draining batteries. Furthermore, transmitting and processing video data on wearable devices is resource-intensive. Many sensing systems (Inertial Measurement Units (IMUs) (Tong et al., 2020; Zhang et al., 2022) or Electromyography (EMG) (Zhang and Amft, 2017, 2018) ) only capture information around the instrumentation position, which is insufficient for activities involving multiple body parts (e.g., face and hands) unless multiple devices are worn on the body, which is less preferred by many users. Microphone-recorded passive sounds (Iravantchi et al., 2021; Mollyn et al., 2022; Laput et al., 2018) have shown promising performance in tracking human activities in the wild. However, it does not work effectively for activities lacking distinctive auditory cues. Therefore, achieving accurate recognition of each nuanced activity often necessitates a new set of customization on wearables with carefully selected sensor combinations, a solution that may not generalize well across different activities. Unlike the computer vision community, which primarily relies on cameras as the sole type of sensors to record data, facilitating the establishment of various types of benchmark datasets, researchers in the wearable and Ubicomp community face challenges in collecting data to track fine-grained human behaviors.
@@
To tackle this challenge, we utilized an empirical observation of human activities that a significant portion of human activities involve both facial movements and upper body movements. If a single sensing system on eyeglasses can simultaneously capture high-quality facial movements and upper body poses, this captured information can potentially be learned by a customized deep learning framework to distinguish a wide set of human activities. However, designing such a wearable system that can accurately recognize various types of human activities with high resolution is highly challenging:
@@
  • Demonstrating the efficacy of the system with high resolution requires recording and labeling ground truth data at the second level, a challenging and time-consuming task that many prior works overlooked.
Recent research has demonstrated the feasibility of integrating active acoustic sensing into eyeglasses using commodity microphones and speakers. By analyzing the reflected acoustic waves from the eyes, face, or upper body, it becomes possible to infer gaze (Li et al., 2024c), facial expression (Zhang et al., 2023b; Li et al., 2024a), or upper body pose (Mahmud et al., 2023; Jin et al., 2021) respectively. This development underscores the potential of leveraging a single sensing modality—active acoustic sensing on eyeglasses—to simultaneously capture both facial movements and upper body poses, which can be used for activity recognition. In this paper, we aim to answer the research question:
@@
Given the small size, relatively low power consumption, low cost, and ubiquity of microphones and speakers, such eyeglasses capable of tracking diverse human activities would significantly reduce the barriers to activity monitoring, a critical task within the Ubicomp community. Instead of relying on multiple devices for different activities, this all-in-one system could effectively track and differentiate between multiple activities simultaneously. However, designing such an active-acoustic sensing system and proving its efficacy for everyday activity recognition presents significant challenges:
@@
To address this research question, we developed ActSonic, a self-supervised and low-power activity recognition system integrated into eyeglasses, utilizing active acoustic sensing. It is the first system to demonstrate the feasibility of recognizing 27 types of everyday activities without collecting training data from new users. Thanks to the low-power nature of acoustic sensors, ActSonic can operate for over 21 hours with a battery capacity equivalent to that of Google Glass (570 mAh).
We developed ActSonic by attaching a pair of miniature, low-power, off-the-shelf microphones and speakers to the hinges of glasses, respectively. The sensing system emits inaudible ultrasonic waves to create an acoustic aura around the body. Based on the shape and position of various body parts, the acoustic signals are reflected with unique and fused patterns captured by the microphone. We developed a customized self-supervised deep learning framework to interpret the reflected signals at areas of interest and their temporal patterns, which are presented with complex multipath echoes and include rich information about movements on both the face and upper body, to infer the performed activities.
@@
  • The first demonstration of creating and utilizing inaudible acoustic waves around the body for fine-grained everyday activity recognition.
@@
  • Compared to many previous works relying on multiple sensors or instrumentations, ActSonic accurately recognizes 27 activities with a single device at a high time resolution of one second.
@@ 2.1. IMU-based Human Activity Recognition
Inertial Measurement Units (IMUs) in commodity smartwatches, phones, and other wearables have garnered considerable interest in detecting human activities over a significant period. These inertial sensors include accelerometers, gyroscopes, magnetometers, heart rate sensors, etc. Early research into IMU-based activity recognition relied on hand-crafted features (Kwon et al., 2018; Qian et al., 2019; Bao and Intille, 2004; Haresamudram et al., 2019) generated from sensor readings. These methods (Kwapisz et al., 2011) were initially limited to recognizing coarse human locomotion activities such as walking, running, sitting, etc. With the advent of end-to-end deep learning methods (Guan and Plötz, 2017; Hammerla et al., 2016; Ordóñez and Roggen, 2016; Murahari and Plötz, 2018; Mahmud et al., 2020) to extract feature representations from time-domain sensor readings, IMU-based systems demonstrated an extended capability to recognize more fine-grained actions such as apparatus usage (Laput and Harrison, 2019), body gestures (Laput et al., 2016), etc. These deep learning architectures, incorporating convolutions, recurrence, and transformers (Ma et al., 2019; Yao et al., 2019; Tonmoy et al., 2021), led to the detection of activities with high fidelity and low error rates. Self-supervision for IMU-based activity recognition is particularly effective in scenarios with small labeled datasets and exhibits significant performance improvement through the utilization of large unlabeled datasets. The pre-training tasks for these self-supervised models are designed as masked window reconstruction (Haresamudram et al., 2020), signal correspondence learning (Saeed et al., 2020), and contrastive predictive coding (Haresamudram et al., 2021) of temporal signals. Despite the success of IMU-based systems in detecting certain fine-grained activities, their capacity is limited due to the low spatial resolution (Tong et al., 2020) of the sensing modalities. Therefore, capturing a wide range of activities with a single placement of IMU remains challenging for wearables.
@@ 2.3. Acoustic Sensing-based Human Activity Recognition
The aforementioned vision-based and multimodal sensing approaches for detecting human activities offer higher spatiotemporal resolution, leading to lower recognition errors. However, the usage of multiple sensors raises concerns about increased power consumption and privacy invasion. Recently, researchers have leveraged the pervasiveness of microphones in off-the-shelf commodity devices to recognize human activities (Wu et al., 2020; Liang and Thomaz, 2019; Laput et al., 2018; Lane et al., 2015). These acoustic sensing systems utilize passively sensed audio within the audible frequency range ($20$ Hz to $16$ KHz). Although passively sensed environmental sound provides discriminative information required to infer certain activities that generate sound, it raises serious privacy concerns since it may record personal conversations. In response to this, recent works (Mollyn et al., 2022; Chen et al., 2008) adopt subsampling and other preprocessing techniques to make the audio unintelligible and then recognize activities based on that. Additionally, passively sensed audio from inaudible frequency ranges (infrasonic and ultrasonic) has been utilized to recognize daily activities (Kijima et al., 2018; Iravantchi et al., 2021, 2019). While these techniques offer better preservation of user privacy, activity recognition on these systems relies on the assumption that activities will generate environmental sound that can be modeled by the system. However, many activities in daily living involve body-limb movement and do not necessarily generate distinctive sounds. Active acoustic sensing-based approaches emit high-frequency sound waves and leverage the reflected sounds to capture fine-grained facial movements (Li et al., 2022; Zhang et al., 2023a; Sun et al., 2023; Li et al., 2024b), hand poses (Lee et al., 2024), sign language gesture (Jin et al., 2021, 2023), body pose (Mahmud et al., 2023), gaze (Li et al., 2024c)and vital signs (Wang et al., 2018; Nandakumar et al., 2015).
@@ 3. Design and Implementation of Sensing System
Our goal is to develop a wearable activity recognition platform integrated into glasses, enabling the identification of a wide range of everyday activities based on the positions and movements of different body parts, as captured by active acoustic sensing technology. While prior research (Mahmud et al., 2023; Li et al., 2022; Zhang et al., 2023a; Li et al., 2024a) has underscored the efficacy of active acoustic sensing in detecting facial changes and upper body poses, no existing system has seamlessly integrated this sensing modality for the comprehensive task of classifying diverse daily activities in real-world scenarios. Given that most daily activities entail movements across various body parts, particularly the upper body, head, and face, our primary focus with ActSonic is to capture and differentiate these activities based on the position and movement of these body areas, as detected by our sensing system. ActSonic accomplishes this with a singular device, facilitating concurrent information capture. This section provides an overview of the active acoustic sensing setup, the process of feature extraction for activity recognition, and the hardware implementation, with a particular emphasis on the wearable form factor of the system.
@@ 3.1. Configuration of Active Acoustic Signal
The active acoustic sensing system in ActSonic integrates two pairs of ultrasonic transmitters and receivers on eyeglass hinges. Utilizing Cross-correlation based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018) chirps, ActSonic emits ultrasonic signals with linearly modulated frequencies ranging from $18$ to $21.5$ kHz and $21.5$ to $24.5$ kHz for the left and right transmitters, respectively, each with a bandwidth of $B=3$ kHz.
@@
The maximum sensing range of ActSonic can be computed as $R_{\text{max}}=N\cdot\delta R=600\cdot 0.343\text{ cm}=205.8\text{ cm}\approx 2\text{ m}$. This combination of high sensing resolution ($0.343$ cm) and extensive sensing range enables us to capture both subtle skin deformations on the face and monitor the pose and coarse movement of the upper body region effectively. Additionally, it’s important to note that the sensing resolution mentioned above is calculated for range-finding applications, as demonstrated in (Wang et al., 2018), which differs from our approach. This calculation is provided as a reference to facilitate readers’ understanding of the sensing principle underlying our proposed system.
@@ 3.2. Computation of Echo Profile and Acoustic Flow
The cross-correlation matches the sweep period of the emitted FMCW ultrasonic wave. It generates an echo frame, represented as a $(600\times 1)$ column vector in ActSonic. Stacking these frames creates the echo profile (Nandakumar et al., 2016; Wang et al., 2018; Li et al., 2022; Zhang et al., 2023a), where brighter pixels indicate strong reflections at specific distances. With two transmitter-receiver pairs, ActSonic accounts for four transmission paths. To elaborate, if we consider the left and right transmitter-receiver pair as $(T_{left},R_{left})$ and $(T_{right},R_{right})$ respectively, then the paths will be $T_{left}\rightarrow R_{left}$, $T_{left}\rightarrow R_{right}$, $T_{right}\rightarrow R_{left}$, and $T_{right}\rightarrow R_{right}$. In ActSonic, we stack the outputs of cross-correlation of these four paths as four channels of the echo profile. Note that the computation of one channel in the echo profile is shown in Figure 2.
@@ 3.3. Hardware Implementation and Wearable Form Factor
We assembled the active acoustic sensing system for ActSonic using two OWR-05049T-38D speakers and two ICS-43434 microphones (TDK, [n. d.]), following a design similar to that shown in (Li et al., 2022). Managed by a Teensy 4.1 microcontroller (PJRC, [n. d.]), the setup oversaw FMCW signal transmission and reception. To connect the speakers, microphones, and microcontroller, we developed a custom PCB housing two MAX98357A audio amplifier chips (maxim integrated, [n. d.]). Utilizing the Inter-IC Sound (I2S) interface, ActSonic’s hardware components communicate, with received signals stored on an SD card via the micro-SD interface on the microcontroller.
@@
Initially, our prototype utilized the Teensy 4.1 microcontroller, powered by an ARM M7 core and exhibiting higher power consumption due to its characteristics. To highlight the power efficiency of our acoustic sensing, we designed a low-power variant featuring an nRF52840 microcontroller (Semiconductors, [n. d.]) (depicted in Figure 3(d)), based on a low-power ARM M4 core. This variant includes two MAX98357A audio amplifiers, similar to the original setup, alongside power management modules and the SGW1110 (SGWireless, [n. d.]) module. A 32 GB SanDisk Extreme microSD card (Amazon.com, [n. d.]) handles storage, with optimized firmware minimizing SD card accesses for swift operations.

### v3

# ActSonic: Recognizing Everyday Activities from Inaudible Acoustic Wave Around the Body

@@ Abstract.
We present ActSonic, an intelligent, low-power active acoustic sensing system integrated into eyeglasses that can recognize 27 different everyday activities (e.g., eating, drinking, toothbrushing) from inaudible acoustic waves around the body. It requires only a pair of miniature speakers and microphones mounted on each hinge of the eyeglasses to emit ultrasonic waves, creating an acoustic aura around the body. The acoustic signals are reflected based on the position and motion of various body parts, captured by the microphones, and analyzed by a customized self-supervised deep learning framework to infer the performed activities on a remote device such as a mobile phone or cloud server. ActSonic was evaluated in user studies with 19 participants across 19 households to track its efficacy in everyday activity recognition. Without requiring any training data from new users (leave-one-participant-out evaluation), ActSonic detected 27 activities, achieving an average F1-score of 86.6% in fully unconstrained scenarios and 93.4% in prompted settings at participants’ homes.
Acoustic Sensing; Activity Recognition, Self-supervised Leaning
copyright: acmlicensedjournal: IMWUTjournalyear: 2024journalvolume: 8journalnumber: 4article: 183publicationmonth: 12doi: 10.1145/3699752ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. INTRODUCTION
Acquiring precise body pose data from various body parts is essential for accurately distinguishing different human activities. However, achieving this with everyday wearable devices (e.g., eyeglasses) has presented persistent challenges: wearable cameras (Price et al., 2022; Liu et al., 2023) consume significant power, rapidly draining batteries. Furthermore, transmitting and processing video data on wearable devices is resource-intensive. Many sensing systems (Inertial Measurement Units (IMUs) (Tong et al., 2020; Zhang et al., 2022) or Electromyography (EMG) (Zhang and Amft, 2017, 2018)) only capture information around the instrumentation position, which is insufficient for activities involving multiple body parts (e.g., face and hands) unless multiple devices are worn on the body, which is less preferred by many users. Microphone-recorded passive sounds (Iravantchi et al., 2021; Mollyn et al., 2022; Laput et al., 2018) have shown promising performance in tracking human activities in the wild. However, it does not work effectively for activities lacking distinctive auditory cues. Therefore, achieving accurate recognition of each nuanced activity often necessitates a new set of customizations on wearables with carefully selected sensor combinations, a solution that may not be generalized well across different activities. Unlike the computer vision community, which primarily relies on cameras as the sole type of sensor to record data, facilitating the establishment of various benchmark datasets, researchers in the wearable and ubiquitous computing community face challenges in collecting data to track fine-grained human behaviors.
@@
To tackle this challenge, we utilized the empirical observation that many human activities involve both facial and upper-body movements. By using a single sensing system on eyeglasses to simultaneously capture high-quality facial movements and upper body poses, this information can potentially be processed by a customized deep learning framework to distinguish a wide range of human activities. However, designing such a wearable system that accurately recognizes various human activities with high resolution is highly challenging:
@@
  • Demonstrating the efficacy of the system with high temporal resolution requires recording and labeling ground truth data at the second level, a challenging and time-consuming task that many prior works overlooked.
Recent research has demonstrated the feasibility of incorporating active acoustic sensing into eyeglasses using commodity microphones and speakers. By analyzing the reflected acoustic waves from the eyes, face, or upper body, it becomes possible to infer gaze (Li et al., 2024c), facial expression (Zhang et al., 2023b; Li et al., 2024a), or upper body pose (Mahmud et al., 2023; Jin et al., 2021; Fan et al., 2023) respectively. This development underscores the potential of leveraging a single sensing modality—active acoustic sensing on eyeglasses—to simultaneously capture both facial movements and upper body poses, which can be used for activity recognition. In this paper, we aim to answer the research question:
@@
Given the small size, relatively low power consumption, low cost, and ubiquity of microphones and speakers, such eyeglasses capable of tracking diverse human activities would significantly reduce the barriers to activity monitoring, a critical task within the ubiquitous computing community. Instead of relying on multiple devices for different activities, this all-in-one system could effectively track and differentiate between multiple activities simultaneously. However, designing such an active-acoustic sensing system and proving its efficacy for everyday activity recognition presents significant challenges:
@@
To address this research question, we developed ActSonic, a self-supervised and low-power activity recognition system placed on eyeglasses form factor, utilizing active acoustic sensing. It is the first system to demonstrate the feasibility of recognizing 27 types of everyday activities without collecting training data from new users. Thanks to the low-power nature of acoustic sensors, ActSonic sensing system for data acquisition can operate for over 21 hours with a battery capacity equivalent to that of Google Glass (570 mAh).
We developed ActSonic by attaching a pair of miniature, low-power, off-the-shelf microphones and speakers to the hinges of glasses. The sensing system emits inaudible ultrasonic waves to create an acoustic aura around the body. Based on the shape and position of various body parts, the acoustic signals are reflected with unique and fused patterns captured by the microphone. We designed a customized self-supervised deep learning framework, running on a cloud server or Android mobile phone, to interpret the reflected signals and their temporal patterns. These signals, presented with complex multipath echoes, include rich information about movements on both the face and upper body, allowing us to infer the performed activities.
@@
  • The first demonstration of creating and utilizing inaudible acoustic waves around the body on eyeglasses form factor for fine-grained everyday activity recognition.
@@
  • Compared to many previous works relying on multiple sensors or instrumentations, ActSonic accurately recognizes 27 activities with a single sensing module at a high time resolution of one second.
@@ 2.1. IMU-based Human Activity Recognition
Inertial Measurement Units (IMUs) in commodity smartwatches, phones, and other wearables have garnered considerable interest in detecting human activities over a significant period. These inertial sensors include accelerometers, gyroscopes, magnetometers, etc. Early research into IMU-based activity recognition relied on hand-crafted features (Kwon et al., 2018; Qian et al., 2019; Bao and Intille, 2004; Haresamudram et al., 2019) generated from sensor readings. These methods (Kwapisz et al., 2011) were initially limited to recognizing coarse human locomotion activities such as walking, running, sitting, etc. With the advent of end-to-end deep learning methods (Guan and Plötz, 2017; Hammerla et al., 2016; Ordóñez and Roggen, 2016; Murahari and Plötz, 2018; Mahmud et al., 2020) to extract feature representations from time-domain sensor readings, IMU-based systems demonstrated an extended capability to recognize more fine-grained actions such as apparatus usage (Laput and Harrison, 2019), body gestures (Laput et al., 2016), etc. These deep learning architectures, incorporating convolutions, recurrence, and transformers (Ma et al., 2019; Yao et al., 2019; Tonmoy et al., 2021), led to the detection of activities with high fidelity and low error rates. Self-supervision for IMU-based activity recognition is particularly effective in scenarios with small labeled datasets and exhibits significant performance improvement through the utilization of large unlabeled datasets. The pre-training tasks for these self-supervised models are designed as masked window reconstruction (Haresamudram et al., 2020), signal correspondence learning (Saeed et al., 2020), and contrastive predictive coding (Haresamudram et al., 2021) of temporal signals. Despite the success of IMU-based systems in detecting certain fine-grained activities, their capacity is limited due to the low spatial resolution (Tong et al., 2020) of the sensing modalities. Therefore, capturing a wide range of activities with a single placement of IMU remains challenging for wearables.
@@ 2.3. Acoustic Sensing-based Human Activity Recognition
The aforementioned vision-based and multimodal sensing approaches for detecting human activities offer higher spatiotemporal resolution, leading to lower recognition errors. However, the usage of multiple sensors raises concerns about increased power consumption and privacy invasion. Recently, researchers have leveraged the pervasiveness of microphones in off-the-shelf commodity devices to recognize human activities (Wu et al., 2020; Liang and Thomaz, 2019; Laput et al., 2018; Lane et al., 2015). These acoustic sensing systems utilize passively sensed audio within the audible frequency range ($20$ Hz to $16$ KHz). Although passively sensed environmental sound provides discriminative information required to infer certain activities that generate sound, it raises serious privacy concerns since it may record personal conversations. In response to this, existing works (Mollyn et al., 2022; Chen et al., 2008; Liang et al., 2020) adopt subsampling and other preprocessing techniques to make the audio unintelligible before activity recognition. Additionally, passively sensed audio from inaudible frequency ranges (infrasonic and ultrasonic) has been utilized to recognize daily activities (Kijima et al., 2018; Iravantchi et al., 2021, 2019). While these techniques offer better preservation of user privacy, activity recognition on these systems relies on the assumption that activities will generate environmental sound that can be modeled by the system. However, many activities in daily living involve body-limb movement and do not necessarily generate distinctive sounds. Active acoustic sensing-based approaches emit high-frequency sound waves and leverage the reflected sounds to capture fine-grained facial movements (Li et al., 2022b; Zhang et al., 2023a; Sun et al., 2023; Li et al., 2024a), hand poses (Lee et al., 2024; Nandakumar et al., 2016; Zhang et al., 2018; Yokota and Hashida, 2016), sign language gesture (Jin et al., 2021, 2023), body pose (Mahmud et al., 2023), activities (Li et al., 2022a), gaze (Li et al., 2024c), vital signs (Wang et al., 2018; Nandakumar et al., 2015), fitness monitoring (Xie et al., 2021), and robot manipulation (Liu et al., 2022; Lu and Culbertson, 2023).
@@ 3. DESIGN AND IMPLEMENTATION OF SENSING SYSTEM
Our goal is to develop a wearable activity recognition platform integrated into glasses, enabling the identification of a wide range of everyday activities based on the positions and movements of different body parts, as captured by active acoustic sensing technology. While prior research (Mahmud et al., 2023; Li et al., 2022b; Zhang et al., 2023a; Li et al., 2024b; Jin et al., 2021; Nandakumar et al., 2016) has underscored the efficacy of active acoustic sensing in detecting facial changes and upper body poses, no existing system has seamlessly integrated this sensing modality for the comprehensive task of classifying diverse daily activities in real-world scenarios. Given that most daily activities entail movements across various body parts, particularly the upper body, head, and face, our primary focus with ActSonic is to capture and differentiate these activities based on the position and movement of these body areas, as detected by our sensing system. ActSonic accomplishes this with a singular sensor placement, facilitating concurrent information capture. This section provides an overview of the active acoustic sensing setup, the process of feature extraction for activity recognition, and the hardware implementation, with a particular emphasis on the wearable form factor of the system.
@@ 3.1. Configuration of Active Acoustic Signal
The active acoustic sensing system in ActSonic incorporates two pairs of ultrasonic transmitters and receivers on the eyeglass hinges. Utilizing Cross-correlation-based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018) chirps, ActSonic emits ultrasonic signals with linearly modulated frequencies ranging from $18.0$ to $21.0$ kHz and $21.5$ to $24.5$ kHz simultaneously for the left and right transmitters, respectively, each with a bandwidth of $B=3$ kHz.
@@
The maximum sensing range of ActSonic can be computed as $R_{\text{max}}=N\cdot\delta R=600\cdot 0.343\text{ cm}=205.8\text{ cm}\approx 2\text{ m}$. This combination of high sensing resolution ($0.343$ cm) and extensive sensing range enables us to capture both subtle skin deformations on the face and monitor the pose and coarse movement of the upper body region effectively. Additionally, it is important to note that the sensing resolution mentioned above is calculated for range-finding applications, as demonstrated in (Wang et al., 2018), which differs from our approach. This calculation is provided as a reference to facilitate readers’ understanding of the sensing principle underlying our proposed system.
@@ 3.2. Computation of Echo Profile and Acoustic Flow
The cross-correlation matches the sweep period of the emitted FMCW ultrasonic wave. It generates an echo frame, represented as a $(600\times 1)$ column vector in ActSonic. Stacking these frames creates the echo profile (Nandakumar et al., 2016; Wang et al., 2018; Li et al., 2022b; Zhang et al., 2023a), where brighter pixels indicate strong reflections at specific distances. With two transmitter-receiver pairs, ActSonic accounts for four transmission paths. To elaborate, if we consider the left and right transmitter-receiver pair as $(T_{left},R_{left})$ and $(T_{right},R_{right})$ respectively, then the paths will be $T_{left}\rightarrow R_{left}$, $T_{left}\rightarrow R_{right}$, $T_{right}\rightarrow R_{left}$, and $T_{right}\rightarrow R_{right}$. In ActSonic, we stack the outputs of cross-correlation of these four paths as four channels of the echo profile. Note that the computation of one channel in the echo profile is shown in Figure 2.
@@ 3.3. Hardware Implementation and Wearable Form Factor
We assembled the active acoustic sensing system for ActSonic using two OWR-05049T-38D speakers and two ICS-43434 microphones (TDK, [n. d.]), following a design similar to that shown in (Li et al., 2022b). Managed by a Teensy 4.1 microcontroller (PJRC, [n. d.]), the setup oversaw FMCW signal transmission and reception. To connect the speakers, microphones, and microcontroller, we developed a custom PCB housing two MAX98357A audio amplifier chips (maxim integrated, [n. d.]). Utilizing the Inter-IC Sound (I2S) interface, ActSonic’s hardware components communicate, with received signals stored on an SD card via the micro-SD interface on the microcontroller.
@@
Initially, our prototype utilized the Teensy 4.1 (PJRC, [n. d.]) microcontroller, powered by an ARM M7 core, which exhibited higher power consumption due to its characteristics. This Teensy 4.1 (PJRC, [n. d.])-based prototype was used to collect data in the user studies intended to evaluate the system. To highlight the power efficiency of our acoustic sensing, we designed a low-power variant featuring an nRF52840 microcontroller (Semiconductors, [n. d.]), based on a low-power ARM M4 core. This variant includes two MAX98357A audio amplifiers, similar to the original setup, alongside power management modules and the SGW1110 (SGWireless, [n. d.]) module. A 32 GB SanDisk Extreme microSD card (Amazon.com, [n. d.]) handles storage, with optimized firmware minimizing SD card accesses for swift operations.